BP13 testing at LFBG, and with AO at RFBG, plus LBC Pointing - 11 October 2011 UT

Observer(s): J. Hill (LBTO) plus Alfio, Lorenzo, Carmelo for first half
Telescope Operator: S. Allanson (LBTO)
Telescope Support: B. Bingham (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

We got to verify TCS Build BP13 at both LFBG and LDG since the SX M2 UMAC got fixed. Thus, testing of BP13 is tolerably complete, and it is deployed for science.

We took about an hour of closed loop AO data on the DX side. And we found a bug in the logic of the (enhanced since BP11) RunAO command that needs to be patched in IIF tomorrow.

We found out that we didn't see Jupiter in AGW2 last night because the detector was completely saturated. So non-sidereal track mode (open loop) works in all cases. Non-sidereal TRACK mode is also tested and ready for science.

We verified the co-pointing and improved focus from Andrew's updated collimation models. (not bad for a guy sitting on a beach!) Then we took binocular pointing data (58 stars) using those new models.

We finished up the night with another hour of non-sidereal observations of Comet Hill. We tested two different observing strategies with and without dithering. A fast OB without guiding hung for 20 minutes at the end of the night, apparently from an RPC error on the Trackers.

Preparations

SX M2 UMAC is apparently repaired with the spare M3 UMAC CPU. This delayed our startup about 45 min.

Bruce also repaired the lock out problem with the LFBG rotator.

Details

Open

01:10 Open to mostly clear sky, with clouds on western horizon. T=5degC Wind 8-10 m/sec. Doug is still filling PISCES.

IRTC (PISCES) @ RFBG

01:50 Going to a pointing star at RFBG

02:00 Initial XY translation of DX M2 Global to correct pointing X=1.050 Y=1.050 Z=0.738 RX=78.3 RY=336.4

02:03 preset to BS9183 star to collimate. Seeing is about 0.7- 0.8. DX M2 ACTIVE X=-0.861 Y=-2.057 Z=-0.096 RX=214.3 RY=-89.7 @EL=62

02:10 preset to offaxis on same BS star.

Team AO is turning on PISCES. We didn't see starlight, so we switched to LFBG.

LUCI (absent) @ LFBG

02:27 preset to M5 star, then to BS9183 star. The image was already well collimated and focussed, but no starlight on WFS. Star overlay was turned on! It was a fake well collimated star.

02:37 ACQUIRE preset to BS9183 star. IE=73 CA=45. Now we have some healthy astigmatism - frighteningly the same as DX had.

02:43 Steve sends various absolute and relative offsets. They work OK.

02:50 Off-axis guide star off the Y-axis

02:53 preset with included hotspot and DETXY offset - failed
Tue Oct 11 02:54:06.223 2011  Preset alert for left LUCIFER mode ACTIVE: guidestar list successfully changed but probe couldn''t be moved to position.
Tue Oct 11 02:54:06.236 2011  Preset command failed for left LUCIFER mode ACTIVE with guidestars, offset, hotspot, Sound=presetFailed.wav

02:55 stop/startAGW, restart GCSL

03:00 preset with included hotspot and DETXY offset - succeeded

BP13 has passed its testing at monocular LFBG for LUCI1.

PISCES @ RFBG

iif_register

03:04 Preset to AO814 with IDL "observe"

default PA of 90 was bad for LN2.

03:07 resend preset with PA=-180, but now guidestar 7 is out of range.

03:11 resend preset with PA=-180, gs=12 - failed for lack of a star. But we were just observing here an hour ago?

03:13 Preset ACQUIRE to ACT star with IRAF. Recentering pointing with XY translation. X=1.350 Y=0.850.

03:18 Changed guidecam_binning 2 for AIP_IRTC_R.cfg

03:21 Preset to AO814 with IDL "observe" PA=150 gs=34 - failed at the issueing of RunAO
Tue Oct 11 03:23:48.829 2011  RunAO command rejected for right IRTC because Invalid type

03:27 Steve now resends a similar preset from IRAF.
Tue Oct 11 03:28:05.860 2011  RunAO command rejected for right IRTC because Invalid type

Alfio figures out that repoint type was true/false in BP11, but it now enumerated in BP13. So the same problem exists in IDL and IRAF.

03:37 Preset to AO814 with IDL "observe" PA=150 gs=34 (now with CHECK=NORMAL specified for RunAO), but that still failed.

Modify the call to iif_runao in observe.pro to force Check=NORMAL. (sent email to Doug for the permanent fix)

03:41 Preset to AO814 with IDL "observe" PA=150 gs=34 - but still failed. We missed an ELSE the first time.

Modify irc_runao to default to NORMAL rather than true.

03:44 Preset to AO814 with IDL "observe" PA=150 gs=34 - RunAO succeeds in getting to AOS, but then......
RunAO command started from right IRTC, type NORMAL, leftRePoint 0, rightRePoint 0
Tue Oct 11 03:48:04.911 2011  RunAO alert for right IRTC: AOARB:Illegal command for state Operational (AcquireRefAO)
Tue Oct 11 03:48:04.949 2011  RunAO alert for right IRTC: AcquireRef[17] failed - AOARB:Illegal command for state Operational (AcquireRefAO)
Tue Oct 11 03:48:04.984 2011  RunAO alert for right IRTC: AcquireRef[135] complete -  dx:0.051 dy:-0.258 slNull: f1:Filter #1 f2:Filter #0 freq:990 mag:8.586 nBins:1 nModes:400 r0:0 snMode:0 strehl:0 ttMod:3
Tue Oct 11 03:48:05.009 2011  RunAO command failed from right IRTC, type NORMAL, leftRePoint 0, rightRePoint 0

It seems that the left side got an AcquireRef 03::46:57 that was rejected (properly) at 03:48:04. ??? Call Chris. There seems to be a logical hole in the RunAO command processing. Chris will send an IIF patch tomorrow after he has reviewed the truth table and the defaults in the light of day. Issue 3725.

03:52:53 both rotators (LFBG and RFBG) faulted together. Call Dan.

Turning on MODS1 agw,calib,hatch

04:12 Authorize MODS @ LDG (another BP13 test) - succeed.

04:15 Authorize back to IRTC @ RFBG

04:18 Preset to AO814 with IDL "observe" PA=90 gs=27 (bad selection as 27 is a triple star)

DX AO Closed Loop Tests

04:20 Preset to AO814 with IDL "observe" PA=90 gs=28 - succeed. Alfio will close the loop manually.

MODS1 @ LDG

05:00 Authorize MODS @ LDG

Open MODS1 hatch

05:02 ACQUIRE preset for MODS - was good except M3 swing arm still deployed. Retract swing arm.

05:15 ACQUIRE preset for MODS - PA mode on-axis (from IRAF Observe.newpoint)

05:17 ACQUIRE preset for MODS - Parallactic mode on-axis

05:21 ACTIVE preset for MODS - star was too far out of focus.

05:23 ACQUIRE preset to adjust focus.

05:29 ACTIVE preset - on-axis - succeeded.

05:32 ACTIVE preset off-Yaxis, included offset, included hotspot, with unwrap - failed for Steve picking wrong PA.

05:37 DETXY offset

05:39 ACQUIRE preset to off-Yaxis star

05:40 NSI TRACK preset to 599 Jupiter.

setGSsfp left 0 0 to center the probe

In /home/telescope/Jupiter/

readGuideCam left -e 100 readGuideCam left -e 5000 - no hint of Jupiter

05:51 ACQUIRE preset to pointing star - pointing is OK (12 arcsec off)

05:53 TRACK preset to the pointing star

setGSsfp left 0 0 to center the probe readGuideCam left -e 100 - we see the star

startAcquire left and center the star IE=-50 CA=+53

05:57 TRACK preset to Jupiter

setGSsfp left

startAcquire left

ptspiral 60 arcsec (finds a glow on the first step)

setFilter left 4 (ND1.0)

Jupiter is still frying the guide camera with ND1. Remember that MODS guider has bigger pixels on sky.

Steve has a theory that a fully centered Jupiter uniformly saturates the camera and it looks blank.

Steve breaks the NSI gui by typing "MB: Io (JI)". It doesn't like the capital MB.

06:14 Preset to moon Io "501" and we see two moons. Non-sidereal TRACK presets now work for LBC and Gregorian foci.

LBC CoPointing Verification

06:22 Deploy swing arms for LBCs

06:26 Turn on LBCs

06:28 Install these two new files from Andrew in Kino Bay for LBC collimation:
DXPMLBCCollimation_Andrew_Mod_1_10th_Oct_2011.dat
SXPMLBCCollimation_Andrew_Mod_1_10th_Oct_2011.dat

Initialize both Primary PSF models, and fixed a typo in the SX one.

06:38 25 biases in LBC, lights are off with lots of moon.

06:42 Close down MODS

Scattered cirrus clouds are moving in from the west. The sky background from moonlit clouds is quite high.

06:57 WT10_280

07:03 DOFPIA

07:08 WT10_280 at EL=63.2
   Current values for the Right telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:  -0.8880 -1.2680 -0.3820  -6.0540  31.1080   0.0000 
   M1Temp:  4.77  Ambient:  5.09 degrees C

   Current values for the Left telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:   0.3420  0.7460 -0.3170  42.2740 -62.2420   0.0000 
   M1Temp:  4.88  Ambient:  4.98 degrees C

   Range-Balanced Values :
 SX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 SX  M1:  -0.6934 -0.3318 -0.3170  19.1513 -40.0277   0.0000 

 DX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 DX  M1:   0.1526 -0.2676 -0.3820  15.4082   8.7838   0.0000 

Sending the calculated optics offsets:
SX: dX = -0.009 dY = -0.008 dRX = -0.167 dRY = 0.184
DX: dX = 0.017 dY = -0.070 dRX = -1.500 dRY = -0.366
          Old      Delta      New
       -------   -------   -------
   CE  -30.166   -15.730   -45.897
   CA -130.776    -8.459  -139.235

ptreset and clear instrument offsets

07:18 WT10_340 @ EL=81.2
   Current values for the Right telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:  -0.7790 -1.0380 -1.2650 -14.2810  39.4530   0.0000 
   M1Temp:  4.76  Ambient:  5.07 degrees C

   Current values for the Left telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:   0.2170  0.8400 -1.0870  31.4400 -69.3730   0.0000 
   M1Temp:  4.85  Ambient:  4.98 degrees C

   Range-Balanced Values :
 SX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 SX  M1:  -0.8565 -0.1776 -1.0870   9.6085 -46.3415   0.0000 

 DX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 DX  M1:   0.3069 -0.1078 -1.2650   5.6751  16.1554   0.0000 

Sending the calculated optics offsets:
SX: dX = 0.040 dY = -0.058 dRX = -1.245 dRY = -0.862
DX: dX = -0.023 dY = -0.030 dRX = -0.637 dRY = 0.489
          Old      Delta      New
       -------   -------   -------
   CE  -30.166    -3.675   -33.841
   CA -130.776     6.191  -124.585

ptreset and remove instrument offsets

07:26 WT_316 @ EL=35.4
   Current values for the Right telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:  -1.1420 -1.3190  0.6500  -5.9180  17.0230   0.0000 
   M1Temp:  4.75  Ambient:  4.89 degrees C

   Current values for the Left telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:   0.5840  0.7810  0.7920  42.8930 -48.8310   0.0000 
   M1Temp:  4.83  Ambient:  4.80 degrees C

   Range-Balanced Values :
 SX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 SX  M1:  -0.4688 -0.3375  0.7920  18.8966 -26.2452   0.0000 

 DX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 DX  M1:  -0.0957 -0.2865  0.6500  16.2340  -5.4240   0.0000 

Sending the calculated optics offsets:
SX: dX = -0.245 dY = -0.123 dRX = -2.641 dRY = 5.255
DX: dX = 0.240 dY = 0.037 dRX = 0.792 dRY = -5.142
          Old      Delta      New
       -------   -------   -------
   CE  -30.166     0.790   -29.377
   CA -130.776   -10.363  -141.139

ptreset and clear instrument offsets

Pointing Model Data for CoPointed LBCs

07:30 Go to 90,120 and reset the mount encoders.

07:35 start pointing logging 20111011_sx.log 20111011_dx.log in /home/telescope/PointingLogs/

07:45 ACT0366

Install Andrew's new collimation models with the Z-temperature corrections included:
DXPMLBCCollimation_Andrew_Mod_2_10th_Oct_2011.dat
SXPMLBCCollimation_Andrew_Mod_2_10th_Oct_2011.dat

Initialize the collimation models and cleared active optics on both sides.

07:52 DOFPIA @ EL=68.4

Andrew missed the zeropoint in Z by a little bit.

SX Active: Xact=0 Yact=0 Zact=-0.425 RX=6.576 RY=-4.263
DX Active: Xact=0 Yact=0 Zact=-0.332 RX=5.609 RY=3.863

08:02 WT10_407

08:05 DOFPIA @ EL=35.7

Focus versus elevation error is significantly improved from earlier, and the zeropoints hold.

SX Active: Xact=0 Yact=0 Zact=-0.404 RX=12.106 RY=-2.845
DX Active: Xact=0 Yact=0 Zact=-0.234 RX=8.381 RY=1.901

Update Zglobals: SX=-0.400 DX=-0.300 and install them in the collimation models.

Initialize and Clear Active Optics on both sides.

08:18 ACT0366

08:20 DOFPIA @ EL=70.4 (slightly bothered by clouds)

SX Active: Xact=0 Yact=0 Zact=-0.006 RX=5.785 RY=-5.212
DX Active: Xact=0 Yact=0 Zact=-0.008 RX=4.574 RY=2.970

08:26 ACT0366 @ EL=70.7
   Current values for the Right telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:  -0.8380 -1.1810 -0.6330 -12.8320  33.4970   0.0000 
   M1Temp:  4.46  Ambient:  4.63 degrees C

   Current values for the Left telescope:
   Coll: X_______Y_______Z________RX_______RY_______RZ____
   M1:   0.2870  0.7780 -0.6160  35.3420 -65.7420   0.0000 
   M1Temp:  4.55  Ambient:  4.63 degrees C

   Range-Balanced Values :
 SX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 SX  M1:  -0.7613 -0.2383 -0.6160  13.5383 -43.2514   0.0000 

 DX  Coll: X_______Y_______Z________RX_______RY_______RZ____
 DX  M1:   0.1985 -0.1599 -0.6330   9.0742  11.2608   0.0000 

Sending the calculated optics offsets:
SX: dX = -0.046 dY = -0.108 dRX = -2.321 dRY = 0.992
DX: dX = 0.038 dY = 0.113 dRX = 2.417 dRY = -0.818
          Old      Delta      New
       -------   -------   -------
   CE  -30.166    -4.719   -34.885
   CA -130.776    -0.273  -131.049
These rangebal values kept for the pointing data.

Selection of stars fro
Real pointing data starts here - that's full binocular pointing data

08:30 ACT0366 @ EL=70.8
08:31 WT10_355 1=66 @ EL=69.4
08:33 WT10_399 2=1 @ EL=20.6
Clouds are getting heavier, but we can still see the pointing stars.
08:36 ACT0398 3=39 @ EL= 33.2
08:38 WT10_364 4=78 @ EL=64.3
08:41 WT10_289 5=13 @ EL=23.7
08:44 ACT0187 6=32 @ EL=33.2 (unwrap)
08:50 ACT0244 7=74
08:53 WT10_089 8=9
08:55 ACT0159 9=26
Images are brutal with background over here near the moon.
08:58 ACT0265 10=70
09:00 WT10_216 11=5
09:04 WT10_318 12=21
09:06 ACT0331 13=45
09:08 ACT0286 14=22
09:10 WT10_269 15=23
09:12 ACT0193 17=51
09:15 WT10_314 16=83
09:16 WT10_078 18=8
09:20 WT10_138 19=29
09:23 ACT0280 20=85
09:26 WT10_242 21=57
09:29 ACT0204 22=12
09:31 WT10_291 23=35
09:33 ACT0327 24=87
09:36 WT10_400 25=64
09:39 ACT0403 26=16
09:41 WT10_407 27=17
09:43 ACT0383 30=2
09:46 WT10_404 31=18
09:48 WT10_329 28=81
09:50 WT10_410 32=40
Clouds are all gone and sky is clear again.
09:52 WT10_370 29=52
09:56 WT10_366 33=79
09:59 ACT0382 34=15
10:02 WT10_351 35=37
10:04 ACT0364 36=61
10:07 WT10_340 37=77
10:09 WT10_295 38=76
10:11 ACT0225 39=33
10:14 WT10_153 40=11
10:16 WT10_155 41=31
10:18 WT10_210 42=54
10:20 ACT0229 43=73
10:22 ACT0176 44=52
10:24 WT10_142 45=27
10:26 ACT0127 46=7
10:27 ACT0180 47=25
10:29 WT10_269 48=48
10:31 ACT0301 49=69
10:33 ACT0303 50=46
10:35 WT10_273 51=4
10:37 ACT0361 52=20
10:39 WT10_322 53=3
10:41 ACT0359 54=44
10:43 WT10_346 55=67
10:45 WT10_286 58=47
10:48 ACT0282 images 105012 and 105007

End of the pointing data

10:51 DOFPIA @ EL=77.7

10:57 Pointing logging stopped

SX Active: Xact=0 Yact=0 Zact=0.118 RX=4.898 RY=-1.151
DX Active: Xact=0 Yact=0 Zact=0.119 RX=3.997 RY=3.186

Non-sidereal comet imaging

10:58 OSUmonitorN2403 for Comet 176P LINEAR. This has 20000 ADU background in 120sec, so we may have to go hunting for brighter comets. FWHM is about 1 arcsec. Note that Target Name and Target Commanded coords are wrong in the headers of these non-sidereal observations since they are read from the LBC OB file. (Maybe LBC was guiding.)

11:14 Wierd stuff happened when I played the ACT0359 preset after cancelling N2403 played mistakenly. Like many rapid fire commands to focus the mirrors.

11:20 DOFPIA @ EL=60

11:25 OSUmonitorN2403 for Comet Hill C/2010 G2 (What can I say, it's my favorite!) Oops, I mistakenly left guiding turned on. (It wasn't what I intended to do, but it could serve as a test of sidereal guiding with a non-sidereal LBC target.)

Cancel preset and queue the comet again. This first batch is straight tracking on the comet, so the comet is fixed on the CCD and the stars appear to move.

11:29 OSUmonitorN2403 for Comet Hill C/2010 G2 without guiding.

11:37 OSUmonitorN2403 for Comet Hill C/2010 G2 without guiding.

11:46 OSUmonitorN2403 for Comet Hill C/2010 G2 without guiding.

11:57 fast OB with 60 sec exposures and 20 dither positions without guiding

~12:20 The fast OB running hangs for ~15 minutes after 18 of 40 exposures. why? Both channels are in preset exposure save mode. Must be something related to the TRACKERS RPC error at 12:11:49. See Issue 3728.

12:35 Have to use lbckill/lbcstart to release it.

12:38 Turn off LBC including Housekeeping as Blue will be dismounted.

Close

12:35 Close T=5.4 degC.

Dave's Pointing Data Analysis

From: Dave Thompson 
Subject: new LBC pointing model
Date: Tue, 11 Oct 2011 20:22:05 -0700

        I made a new pointing model from the LBCR data taken last night.  While there were several images I could not use because I could not fit a WCS to the few stars in the images, there were 43 that went into this model.  I would say we are making progress.  The previous model had several extra terms in it to reach 4" rms while this new one has just the standard terms plus PZZ1 (a linear "tube flexure" term) and we are down to 3.3" rms.  I tried the extra terms from previous models (also LUCI models) but they do not actually drop the rms so I removed them.

        I have copied the new pointing model to /home/telescope/TCS/Configuration/PCS/ with a filename of LBCDX.20111011.ptmod.  I did not update the !PCSInstrument.conf file to make use of this.  LBCB should use the same model (since it is now co-pointed).

-bash-3.2$ more LBCDX.20111011.ptmod
LBT right Pointing Log, 2011 Oct 11
T   43   3.2577    0.000   0.0000               #My additional comments (not in ptmod file)
  IA       -331.1000     7.44134                #(Index in Azimuth = encoder zero point)
  IE        -37.6079     1.39394                #(Index in Elevation = encoder zero point)
  NPAE      +46.9460     8.00938        #(Non-Perpedicularity between AZ and EL)
  CA       -109.5057    10.31899                #(local correction orthogonal to EL direction)
  AW         +6.7723     0.60159                #(Azimuth axis offset from zenith E-W)
  AN        +20.5135     0.58917                #(Azimuth axis offset from zenith N-S)
  PZZ1      +32.8920     1.64615                #(1st order Polynomial in Zenith distance)
END

-- JohnHill - 10 Oct 2011
Topic revision: r7 - 17 Oct 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback